PLEASE HELP ME IF ANYONE CAN..(IF 2X+7=12 THEN 4X-7=
OPTIONS ARE
(A)2
(B)2.5
(C)3
(D)10
(E)13

Answer by ewatrrr(24785)

please do not uses CAPS on this site:
2x + 7 = 12
x = 5/2
And......
4(5/2) - 7 = 3